OpenStudy (anonymous):

Find the function value. cot 30° = ? Can someone explain on how to get the answer ? Thanks!

OpenStudy (anonymous):

cot theta =1/tan theta

OpenStudy (anonymous):

cot 30=1/tan 30

OpenStudy (anonymous):

so if it was cot45 it would be cot45=1/tan45 ?

OpenStudy (anonymous):

its invere function of each other yes

OpenStudy (anonymous):

Oh okay. Well how do you get a fraction ? Like I don't know how to find out what tan30 equals.

OpenStudy (anonymous):

tan theta =sin theta/cos theta tan 30=sin 30/cos 60 wait I will send you trick

hartnn (hartnn):

*tan 30=sin 30/cos 30

OpenStudy (anonymous):